# Finite Distance Corrections to Vacuum Birefringence in Strong Gravitational and Electromagnetic Fields

### v1

@@ II The geometric foundation of photon propagation
In this section, we establish the geometric framework necessary to evaluate the finite-distance birefringence effects induced by Non-Linear Electrodynamics (NLED). Our primary objective is to derive the polarization-dependent effective optical metrics, denoted as $\gamma_{ij}^{\pm}$, by analyzing the interaction between the background spacetime geometry and the non-linear electromagnetic Lagrangian, which provides the rigorous basis for applying the Gauss-Bonnet theorem in subsequent sections.
@@ II.1 The background and the field
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-A(r)dt^{2}+B(r)dr^{2}+r^{2}d\Omega^{2},$ (1)
where the metric functions $A(r)$ and $B(r)$ depend solely on the radial coordinate $r$. The angular element is defined as $d\Omega^{2}\equiv d\theta^{2}+\sin^{2}\theta d\phi^{2}$. Equation (1) serves as the background gravitational field and is assumed to be a solution to the Einstein field equations coupled to an electromagnetic source.
We introduce the electromagnetic field tensor $F_{\mu\nu}$, defined as the antisymmetrized derivative of the gauge potential, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. To preserve the static and spherical symmetry of the background spacetime, the electromagnetic field tensor must share the isometries of the metric $g_{\mu\nu}$. Consequently, the independent non-vanishing field components are restricted to the radial electric field $E(r)=F_{tr}$ and the radial magnetic field $B(r)=F_{\theta\phi}/\sin\theta$.
We formally encapsulate the geometric and field-theoretic inputs of our model in the following definitions:
$\boxed{\,\begin{aligned} g_{\mu\nu}&=\text{diag}\left(-A(r),B(r),r^{2},r^{2}\sin^{2}\theta\right),\\ F&=-\frac{1}{2}\frac{E(r)^{2}}{A(r)B(r)}+\frac{1}{2}\frac{B(r)^{2}}{r^{4}},\quad G=-\frac{E(r)B(r)}{r^{2}\sqrt{A(r)B(r)}}.\end{aligned}\,}$ (2)
We have something to remark about the topology: It is important to emphasize that the assumption of a spherically symmetric radial magnetic field ($B(r)\neq 0$) implies the existence of a magnetic monopole charge $q_{m}$ [Bronnikov:2000vy]. While astrophysical magnetars possess dipole fields that break spherical symmetry, we employ this monopole configuration as a monopole approximation. It allows us to analytically model the effects of extreme magnetic field intensities on photon propagation within the spherically symmetric framework of the Gauss-Bonnet theorem, serving as a foundational step toward more complex geometries.
@@ II.2 The NLED Lagrangian and effective geodesics
Having established the background geometry and the electromagnetic invariants, we now turn to the dynamics of the photon field governed by non-linear electrodynamics (NLED). We postulate a generic action for the electromagnetic sector, defined by the Lagrangian density $\mathcal{L}(F,G)$:
$S_{\text{EM}}=\int d^{4}x\sqrt{-g}\mathcal{L}(F,G).$ (3)
@@
$\nabla_{\mu}\left(\mathcal{L}_{F}F^{\mu\nu}+\mathcal{L}_{G}\tilde{F}^{\mu\nu}\right)=0,$ (4)
where $\mathcal{L}_{F}\equiv\partial\mathcal{L}/\partial F$ and $\mathcal{L}_{G}\equiv\partial\mathcal{L}/\partial G$. Unlike in linear Maxwell theory (where $\mathcal{L}=F$), the coefficients $\mathcal{L}_{F}$ and $\mathcal{L}_{G}$ here are functions of the field strength, inducing self-interaction and non-linear propagation effects.
@@
Linearizing the field equations (4) with respect to the perturbation yields a wave equation for the photon. The non-linearity of the Lagrangian results in the background field acting as an effective medium. The principal part of this linearized equation leads to the Fresnel equation for the wave vector $k_{\mu}$ [Boillat:1970gw]. Critically, the non-trivial dependence of $\mathcal{L}$ on $F$ and $G$ causes the Fresnel equation to factorize into two distinct dispersion relations, corresponding to two polarization modes (denoted by $\pm$). The factorization implies that photons do not traverse the null geodesics of the background metric $g_{\mu\nu}$ (Eq. (1)). Instead, they follow the null geodesics of polarization-dependent effective metrics, denoted $g_{\mu\nu}^{\text{eff},\pm}$.
To explicitly derive these metrics, we utilize the formalism established by Novello -et al. and De Lorenci -et al. [Novello:1999pg, DeLorenci:2000yh]. For a background geometry defined by Eq. (2) with radial fields, the cross-terms in the Fresnel equation vanish, and the effective metric tensor diagonalizes. The contravariant components for the two polarization modes are given by:
$\boxed{\,g_{\text{eff},\pm}^{\mu\nu}=\mathcal{L}_{F}g^{\mu\nu}-4\Lambda_{\pm}F^{\mu}_{\ \lambda}F^{\nu\lambda},\,}$ (5)
@@
$\Lambda_{\pm}=\frac{1}{2}\left(\mathcal{L}_{FF}+\frac{1}{2}g^{2}\mathcal{L}_{GG}\pm\sqrt{\left(\mathcal{L}_{FF}-\frac{1}{2}g^{2}\mathcal{L}_{GG}\right)^{2}+g^{2}\mathcal{L}_{FG}^{2}}\right),$ (6)
where $g$ is a scalar function of the field configuration defined by $G/F$ ratios. In the specific case of purely magnetic or purely electric backgrounds often used in SSS models, these expressions simplify, identifying one mode that propagates along the background light cone (the ordinary ray in linear theory limits) and one that deviates (the extraordinary ray).
@@ II.3 Construction of the dual optical manifolds
The effective metric tensor $g_{\mu\nu}^{\text{eff},\pm}$, defined by the null propagation condition in Eq. (5), inherits the static and spherical symmetries of the background metric $g_{\mu\nu}$ and the field tensor $F_{\mu\nu}$. Consequently, the effective line element for photon propagation can be diagonalized and written in a general SSS form. We introduce the effective metric functions $A_{\pm}(r)$, $B_{\pm}(r)$, and $H_{\pm}(r)$ to describe the geometry seen by the two polarization modes:
$ds_{\text{eff},\pm}^{2}=-A_{\pm}(r)dt^{2}+B_{\pm}(r)dr^{2}+H_{\pm}(r)r^{2}d\Omega^{2},$ (7)
where $H_{\pm}(r)$ accounts for the fact that the angular part of the effective metric may differ from the background $r^{2}$ due to the non-linear interaction terms $\mathcal{L}_{FF}$ and $\mathcal{L}_{GG}$ identified in Eq. (5). Note that in the limit of linear electrodynamics, $A_{\pm}\to A$, $B_{\pm}\to B$, and $H_{\pm}\to 1$. Note that since the effective metric tensor derived from the Fresnel equation (Eq. (5)) is contravariant, the functions appearing in the line element (Eq. (7)) are the inverses of the diagonal components: $A_{\pm}(r)=-(g^{tt}_{\text{eff},\pm})^{-1}$ and $B_{\pm}(r)=(g^{rr}_{\text{eff},\pm})^{-1}$. The function $H_{\pm}(r)$ encapsulates the deformation of the angular sector relative to the radial sector induced by the non-linearities. Explicitly, it is defined by the ratio of the effective metric components normalized by the background:
$H_{\pm}(r)=\frac{g^{\theta\theta}_{\text{eff},\pm}/g^{\theta\theta}}{g^{rr}_{\text{eff},\pm}/g^{rr}},$ (8)
@@
To study the spatial trajectory of light rays, we restrict our analysis to the equatorial plane ($\theta=\pi/2$) without loss of generality. By setting the effective line element $ds_{\text{eff},\pm}^{2}=0$, we can solve for the coordinate time interval $dt$ in terms of the spatial coordinates, which allows the projection of the four-dimensional null geodesics onto three-dimensional spatial curves. The resulting quadratic form defines the optical metric $\gamma_{ij}^{\pm}$ for each polarization mode [Birula_1986]:
$dt^{2}=\gamma_{ij}^{\pm}dx^{i}dx^{j}=\frac{B_{\pm}(r)}{A_{\pm}(r)}dr^{2}+\frac{H_{\pm}(r)r^{2}}{A_{\pm}(r)}d\phi^{2}.$ (9)
Equation (9) defines two distinct three-dimensional Riemannian manifolds, which we denote as $M^{opt}_{+}$ and $M^{opt}_{-}$. These manifolds are the optical spaces in which the photons travel as spatial geodesics. The non-vanishing components of the optical metric tensors $\gamma_{ij}^{\pm}$ on the equatorial plane are explicitly identified as:
$\boxed{\gamma_{rr}^{\pm}(r)=\frac{B_{\pm}(r)}{A_{\pm}(r)},\quad\gamma_{\phi\phi}^{\pm}(r)=\frac{H_{\pm}(r)r^{2}}{A_{\pm}(r)}.}$ (10)
@@ III The finite-distance Gauss-Bonnet formalism
In this section, we formulate the deflection angle $\alpha_{\pm}$ for a finite-distance observer using the topology of the optical manifolds $M^{opt}_{\pm}$ derived in Section II. Our approach generalizes the work of Gibbons and Werner [Gibbons:2008rj], who established that the asymptotic deflection angle can be computed as a global topological invariant, which is specifically the surface integral of the Gaussian curvature over the optical manifold. We extend this correspondence to the finite-distance regime by rigorously defining the geometric angles at the observer and source locations within the curved optical space, allowing us to quantify the birefringence $\Delta\alpha$ as a differential geometric effect arising from the distinct curvatures $K^{+}$ and $K^{-}$ of the polarization-dependent manifolds.
@@ III.1 Gaussian curvature of the optical manifolds
To apply the Gauss-Bonnet theorem, we must first determine the intrinsic curvature of the optical submanifolds defined by the metric in Eq. (10). The optical metric for the equatorial plane is given by the 2-dimensional line element:
$dl_{\pm}^{2}=\gamma_{rr}^{\pm}(r)dr^{2}+\gamma_{\phi\phi}^{\pm}(r)d\phi^{2},$ (11)
@@
$\Gamma^{r}_{rr}=\frac{1}{2\gamma_{rr}^{\pm}}\frac{d\gamma_{rr}^{\pm}}{dr},\quad\Gamma^{r}_{\phi\phi}=-\frac{1}{2\gamma_{rr}^{\pm}}\frac{d\gamma_{\phi\phi}^{\pm}}{dr},\quad\Gamma^{\phi}_{r\phi}=\frac{1}{2\gamma_{\phi\phi}^{\pm}}\frac{d\gamma_{\phi\phi}^{\pm}}{dr}.$ (12)
@@
$K^{\pm}(r)=-\frac{1}{\sqrt{\gamma_{rr}^{\pm}\gamma_{\phi\phi}^{\pm}}}\frac{d}{dr}\left(\frac{1}{\sqrt{\gamma_{rr}^{\pm}}}\frac{d\sqrt{\gamma_{\phi\phi}^{\pm}}}{dr}\right).$ (13)
Substituting the expressions for $\gamma_{rr}^{\pm}$ and $\gamma_{\phi\phi}^{\pm}$ from Eq. (10) into Eq. (13), we derive the explicit form of the Gaussian curvature for each polarization mode. Rather than expanding the derivatives fully, it is physically more insightful to present the curvature in a compact factorized form, which reveals how the non-linear effective metric functions modify the standard geometric curvature:
$\boxed{K^{\pm}(r)=-\frac{A_{\pm}(r)}{r\sqrt{B_{\pm}(r)H_{\pm}(r)}}\frac{d}{dr}\left(\sqrt{\frac{A_{\pm}(r)}{B_{\pm}(r)}}\frac{d}{dr}\left(r\sqrt{\frac{H_{\pm}(r)}{A_{\pm}(r)}}\right).\right)}$ (14)
Equation (14) explicitly demonstrates that the birefringence arises from the non-trivial radial gradients of the ratio $H_{\pm}/A_{\pm}$ and the conformal factor $A_{\pm}/B_{\pm}$. In the weak-field limit where $A_{\pm}\approx 1-2M/r$ and $H_{\pm}\approx 1$, this expression reduces to the standard Schwarzschild optical curvature $K\approx-2M/r^{3}$. However, in the strong-field regime near the photon sphere, the term $H_{\pm}(r)$ (containing $\mathcal{L}_{FF}$ and $\mathcal{L}_{GG}$) introduces significant deviations.
It is crucial to analyze the behavior of $K^{\pm}$ near the photon sphere radius $r_{ph}^{\pm}$. The photon sphere is defined by the singularity of the effective potential, or equivalently, where the optical metric components diverge or the geodesic equation allows for circular orbits. If $K^{\pm}$ possesses a singularity at $r_{ph}^{\pm}$, the Gauss-Bonnet integral must be handled with regularization techniques. For standard NLED models like Euler-Heisenberg, the curvature remains regular outside the horizon, but for models with Born-Infeld type structures, divergences may appear that require restricting the integration domain to $r>r_{ph}^{\pm}$ [Li:2019mqw]. The behavior of the divergence fundamentally distinguishes the NLED optical manifold from the vacuum Schwarzschild case.

### v2

@@ II The geometric foundation of photon propagation
In this section, we establish the geometric framework necessary to evaluate the finite-distance birefringence effects induced by NLED. Our primary objective is to derive the polarization-dependent effective optical metrics, denoted as $\gamma_{ij}^{\pm}$, by analyzing the interaction between the background spacetime geometry and the non-linear electromagnetic Lagrangian, which provides the rigorous basis for applying the Gauss-Bonnet theorem in subsequent sections.
@@ II.1 The background and the field
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-A(r)dt^{2}+B(r)dr^{2}+r^{2}d\Omega^{2},$ (II.1)
where the metric functions $A(r)$ and $B(r)$ depend solely on the radial coordinate $r$. The angular element is defined as $d\Omega^{2}\equiv d\theta^{2}+\sin^{2}\theta d\phi^{2}$. Equation (II.1) serves as the background gravitational field and is assumed to be a solution to the Einstein field equations coupled to an electromagnetic source.
We introduce the electromagnetic field tensor $F_{\mu\nu}$, defined as the antisymmetrized derivative of the gauge potential, $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. To preserve the static and spherical symmetry of the background spacetime, the electromagnetic field tensor must share the isometries of the metric $g_{\mu\nu}$. Consequently, the independent non-vanishing field components are restricted to the radial electric field $E(r)=F_{tr}$ and the radial magnetic field strength $\mathcal{B}(r)\equiv F_{\theta\phi}/\sin\theta$.
We formally encapsulate the geometric and field-theoretic inputs of our model in the following definitions. We adopt metric signature (,+,+,+)(-,+,+,+) and define the dual field strength by $\tilde{F}^{\mu\nu}\equiv\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$, with $\epsilon^{tr\theta\phi}=+1/\sqrt{-g}$. The two electromagnetic invariants are $F\equiv\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ and $G\equiv\frac{1}{4}F_{\mu\nu}\tilde{F}^{\mu\nu}$. For a static, spherically symmetric background with radial fields $E(r)=F_{tr}$ and $\mathcal{B}(r)\equiv F_{\theta\phi}/\sin\theta$, these reduce to:
$g_{\mu\nu}$ $=\mathrm{diag}\!\left(-A(r),\,B(r),\,r^{2},\,r^{2}\sin^{2}\theta\right),$
$F$ $\equiv\frac{1}{4}F_{\mu\nu}F^{\mu\nu}=\frac{1}{2}\left(\frac{\mathcal{B}(r)^{2}}{r^{4}}-\frac{E(r)^{2}}{A(r)B(r)}\right),$
$G$ $\equiv\frac{1}{4}F_{\mu\nu}\tilde{F}^{\mu\nu}=\frac{E(r)\,\mathcal{B}(r)}{r^{2}\sqrt{A(r)B(r)}}.$ (II.2)
It is important to emphasize that the assumption of a spherically symmetric radial magnetic field ($\mathcal{B}(r)\neq 0$) implies the existence of a magnetic monopole charge $q_{m}$ [22]. While astrophysical magnetars possess dipole fields that break spherical symmetry, we employ this monopole configuration as a monopole approximation. It allows us to analytically model the effects of extreme magnetic field intensities on photon propagation within the spherically symmetric framework of the Gauss-Bonnet theorem, serving as a foundational step toward more complex geometries.
@@ II.2 The NLED Lagrangian and effective geodesics
Having established the background geometry and the electromagnetic invariants, we now turn to the dynamics of the photon field governed by NLED. We postulate a generic action for the electromagnetic sector, defined by the Lagrangian density $\mathcal{L}(F,G)$:
$S_{\text{EM}}=\int d^{4}x\sqrt{-g}\mathcal{L}(F,G).$ (II.3)
@@
$\nabla_{\mu}\left(\mathcal{L}_{F}F^{\mu\nu}+\mathcal{L}_{G}\tilde{F}^{\mu\nu}\right)=0,$ (II.4)
where $\mathcal{L}_{F}\equiv\partial\mathcal{L}/\partial F$ and $\mathcal{L}_{G}\equiv\partial\mathcal{L}/\partial G$. Unlike in linear Maxwell theory (where $L=-F=-\tfrac{1}{4}F_{\mu\nu}F^{\mu\nu}$), the coefficients $\mathcal{L}_{F}$ and $\mathcal{L}_{G}$ here are functions of the field strength, inducing self-interaction and non-linear propagation effects.
@@
Linearizing the field equations (II.4) with respect to the perturbation yields a wave equation for the photon. The non-linearity of the Lagrangian results in the background field acting as an effective medium. The principal part of this linearized equation leads to the Fresnel equation for the wave vector $k_{\mu}$ [13]. Critically, the non-trivial dependence of $L$ on $F$ and $G$ generically yields a quartic Fresnel equation that factorizes into the product of two quadratic dispersion relations, corresponding to two polarization modes (denoted by $\pm$). In exceptional models (notably Born-Infeld electrodynamics) the two quadratics coincide so the Fresnel polynomial is a perfect square; in that case there is no vacuum birefringence (a single effective light cone shared by both polarizations) [13, 67, 42]. The factorization implies that photons do not traverse the null geodesics of the background metric $g_{\mu\nu}$ (Eq. (II.1)). Instead, they follow the null geodesics of polarization-dependent effective metrics, denoted $g_{\mu\nu}^{\text{eff},\pm}$.
To explicitly derive these metrics, we utilize the formalism established by Novello et al. and De Lorenci et al. [65, 30]. For a background geometry defined by Eq. (II.2) with radial fields, the cross-terms in the Fresnel equation vanish, and the effective metric tensor diagonalizes. The contravariant components for the two polarization modes are given by:
$g_{\text{eff},\pm}^{\mu\nu}=\mathcal{L}_{F}g^{\mu\nu}-4\Lambda_{\pm}F^{\mu}_{\ \lambda}F^{\nu\lambda},\,$ (II.5)
@@
$\Lambda_{\pm}=\frac{1}{2}\left[L_{FF}+\frac{1}{2}\eta^{2}L_{GG}\pm\sqrt{\left(L_{FF}-\frac{1}{2}\eta^{2}L_{GG}\right)^{2}+\eta^{2}L_{FG}^{2}}\right],$ (II.6)
where $L_{FF}\equiv\partial^{2}L/\partial F^{2}$, $L_{GG}\equiv\partial^{2}L/\partial G^{2}$, and $L_{FG}\equiv\partial^{2}L/(\partial F\,\partial G)$ are evaluated on the background, and where we introduced the dimensionless scalar
$\eta^{2}\equiv 1+\left(\frac{G}{F}\right)^{2}=\frac{F^{2}+G^{2}}{F^{2}}.$ (II.7)
In particular, for purely magnetic or purely electric backgrounds one has $G=0$ and hence $\eta=1$. In the Maxwell limit L=FL=-F, one has $L_{F}=-1$ and $L_{FF}=L_{GG}=L_{FG}=0$, implying $\Lambda_{\pm}=0$ and thus Eq. (II.5) reduces (up to an overall conformal factor) to $g^{\mu\nu}_{\mathrm{eff},\pm}\propto g^{\mu\nu}$, as expected. In the specific case of purely magnetic or purely electric backgrounds often used in SSS models, these expressions simplify, identifying one mode that propagates along the background light cone (the ordinary ray in linear theory limits) and one that deviates (the extraordinary ray).
@@ II.3 Construction of the dual optical manifolds
The effective metric tensor $g_{\mu\nu}^{\text{eff},\pm}$, defined by the null propagation condition in Eq. (II.5), inherits the static and spherical symmetries of the background metric $g_{\mu\nu}$ and the field tensor $F_{\mu\nu}$. Consequently, the effective line element for photon propagation can be diagonalized and written in a general SSS form. We introduce the effective metric functions $A_{\pm}(r)$, $B_{\pm}(r)$, and $H_{\pm}(r)$ to describe the geometry seen by the two polarization modes:
$ds_{\text{eff},\pm}^{2}=-A_{\pm}(r)dt^{2}+B_{\pm}(r)dr^{2}+H_{\pm}(r)r^{2}d\Omega^{2},$ (II.8)
where $H_{\pm}(r)$ accounts for the fact that the angular part of the effective metric may differ from the background $r^{2}$ due to the non-linear interaction terms $\mathcal{L}_{FF}$ and $\mathcal{L}_{GG}$ identified in Eq. (II.5). Note that in the limit of linear electrodynamics, $A_{\pm}\to A$, $B_{\pm}\to B$, and $H_{\pm}\to 1$. Note that since the effective metric tensor derived from the Fresnel equation (Eq. (II.5)) is contravariant, the functions appearing in the line element (Eq. (II.8)) are the inverses of the diagonal components: $A_{\pm}(r)=-(g^{tt}_{\text{eff},\pm})^{-1}$ and $B_{\pm}(r)=(g^{rr}_{\text{eff},\pm})^{-1}$. The function $H_{\pm}(r)$ encapsulates the deformation of the angular sector relative to the radial sector induced by the non-linearities. Explicitly, it is defined by the ratio of the effective metric components normalized by the background:
$H_{\pm}(r)=\frac{g^{\theta\theta}_{\text{eff},\pm}/g^{\theta\theta}}{g^{rr}_{\text{eff},\pm}/g^{rr}},$ (II.9)
@@
To study the spatial trajectory of light rays, we restrict our analysis to the equatorial plane ($\theta=\pi/2$) without loss of generality. By setting the effective line element $ds_{\text{eff},\pm}^{2}=0$, we can solve for the coordinate time interval $dt$ in terms of the spatial coordinates, which allows the projection of the four-dimensional null geodesics onto three-dimensional spatial curves. The resulting quadratic form defines the optical metric $\gamma_{ij}^{\pm}$ for each polarization mode [12]:
$dt^{2}=\gamma_{ij}^{\pm}dx^{i}dx^{j}=\frac{B_{\pm}(r)}{A_{\pm}(r)}dr^{2}+\frac{H_{\pm}(r)r^{2}}{A_{\pm}(r)}d\phi^{2}.$ (II.10)
Equation (II.10) defines two distinct three-dimensional Riemannian manifolds, which we denote as $M^{opt}_{+}$ and $M^{opt}_{-}$. These manifolds are the optical spaces in which the photons travel as spatial geodesics. The non-vanishing components of the optical metric tensors $\gamma_{ij}^{\pm}$ on the equatorial plane are explicitly identified as:
$\gamma_{rr}^{\pm}(r)=\frac{B_{\pm}(r)}{A_{\pm}(r)},\quad\gamma_{\phi\phi}^{\pm}(r)=\frac{H_{\pm}(r)r^{2}}{A_{\pm}(r)}.$ (II.11)
@@ III The finite-distance Gauss-Bonnet formalism
In this section, we formulate the deflection angle $\alpha_{\pm}$ for a finite-distance observer using the topology of the optical manifolds $M^{opt}_{\pm}$ derived in Section II. Our approach generalizes the work of Gibbons and Werner [41, 69, 27], who established that the asymptotic deflection angle can be computed as a global topological invariant, which is specifically the surface integral of the Gaussian curvature over the optical manifold. We extend this correspondence to the finite-distance regime by rigorously defining the geometric angles at the observer and source locations within the curved optical space, allowing us to quantify the birefringence $\Delta\alpha$ as a differential geometric effect arising from the distinct curvatures $K^{+}$ and $K^{-}$ of the polarization-dependent manifolds.
@@ III.1 Gaussian curvature of the optical manifolds
To apply the Gauss-Bonnet theorem, we must first determine the intrinsic curvature of the optical submanifolds defined by the metric in Eq. (II.11). The optical metric for the equatorial plane is given by the 2-dimensional line element [70]:
$dl_{\pm}^{2}=\gamma_{rr}^{\pm}(r)dr^{2}+\gamma_{\phi\phi}^{\pm}(r)d\phi^{2},$ (III.1)
@@
$\Gamma^{r}_{rr}=\frac{1}{2\gamma_{rr}^{\pm}}\frac{d\gamma_{rr}^{\pm}}{dr},\quad\Gamma^{r}_{\phi\phi}=-\frac{1}{2\gamma_{rr}^{\pm}}\frac{d\gamma_{\phi\phi}^{\pm}}{dr},\quad\Gamma^{\phi}_{r\phi}=\frac{1}{2\gamma_{\phi\phi}^{\pm}}\frac{d\gamma_{\phi\phi}^{\pm}}{dr}.$ (III.2)
@@
$K^{\pm}(r)=-\frac{1}{\sqrt{\gamma_{rr}^{\pm}\gamma_{\phi\phi}^{\pm}}}\frac{d}{dr}\left(\frac{1}{\sqrt{\gamma_{rr}^{\pm}}}\frac{d\sqrt{\gamma_{\phi\phi}^{\pm}}}{dr}\right).$ (III.3)
Substituting the expressions for $\gamma_{rr}^{\pm}$ and $\gamma_{\phi\phi}^{\pm}$ from Eq. (II.11) into Eq. (III.3), we derive the explicit form of the Gaussian curvature for each polarization mode. Rather than expanding the derivatives fully, it is physically more insightful to present the curvature in a compact factorized form, which reveals how the non-linear effective metric functions modify the standard geometric curvature:
$K^{\pm}(r)=-\frac{A_{\pm}(r)}{r\sqrt{B_{\pm}(r)H_{\pm}(r)}}\frac{d}{dr}\left(\sqrt{\frac{A_{\pm}(r)}{B_{\pm}(r)}}\frac{d}{dr}\left(r\sqrt{\frac{H_{\pm}(r)}{A_{\pm}(r)}}\right).\right)$ (III.4)
Equation (III.4) explicitly demonstrates that the birefringence arises from the non-trivial radial gradients of the ratio $H_{\pm}/A_{\pm}$ and the conformal factor $A_{\pm}/B_{\pm}$. In the weak-field limit where $A_{\pm}\approx 1-2M/r$ and $H_{\pm}\approx 1$, this expression reduces to the standard Schwarzschild optical curvature $K\approx-2M/r^{3}$. However, in the strong-field regime near the photon sphere, the term $H_{\pm}(r)$ (containing $\mathcal{L}_{FF}$ and $\mathcal{L}_{GG}$) introduces significant deviations.
It is crucial to analyze the behavior of $K^{\pm}$ near the photon sphere radius $r_{ph}^{\pm}$. The photon sphere is defined by the singularity of the effective potential, or equivalently, where the optical metric components diverge or the geodesic equation allows for circular orbits [59]. If $K^{\pm}$ possesses a singularity at $r_{ph}^{\pm}$, the Gauss-Bonnet integral must be handled with regularization techniques. For standard NLED models like Euler-Heisenberg, the curvature remains regular outside the horizon, but for models with Born-Infeld type structures, divergences may appear that require restricting the integration domain to $r>r_{ph}^{\pm}$. The behavior of the divergence fundamentally distinguishes the NLED optical manifold from the vacuum Schwarzschild case. In the remainder of this work we restrict to the weak-deflection regime ($b\gg M$ and $r_{0}\simeq b$ well outside $r_{\rm ph}^{\pm}$), so the near-photon-sphere regularization issues mentioned here do not enter our analytic evaluation; extending the construction to strong deflection is left for future work.